2.50.00 mL of concentrated nitric acid (16M

HN03) is diluted to 1.000 L. What is the

concentration of the new solution?


1
Expert's answer
2022-05-20T09:14:01-0400

Given:

c1 = 16 M

V1 = 50.00 mL = 0.050 L

c2 = unknown

V2 = 1.000 L


Solution:


where:

c1, V1 - the molarity and volume of the concentrated solution

c2, V2 - the molarity and volume of the diluted solution


Thus:

(16 M) × (0.050 L) = c2 × (1.000 L)

c2 = (16 M × 0.050 L) / (1.000 L) = 0.8 M

c2 = 0.8 M


Answer: The concentration of the new solution is 0.8 M